# No Metric to Rule Them All: Toward Principled Evaluations of Graph-Learning Datasets

### v2

@@ D.2.4 Supplementary Extensibility Experiments
$\text{Jaccard}(A,A’)$ ’ ) $=\frac{|A\cap A’|}{|A\cup A’|},\ \text{Asymm}(A’,A)=\frac{|A\cap A’|}{|A’|}\;,$
@@
[…]
@@
Dataset descriptions. QM9 (Wu et al., 2018a) is a molecular dataset of 133 885133885133\,885133 885 graphs. Its name originates from its focus on quantum mechanical properties and the makeup of the dataset: stable, organic molecules with nine or fewer heavy atoms. The PyG version has eleven node features, the first five being a one-hot encoding of the atom type, and the rest representing measures such as the hybridization state and the number of hydrogen atoms. Nineteen regression tasks seek to predict a variety molecular properties. We focus on task 00 (as defined by PyG), regressing the dipole moment, which describes the spatial distribution of electrons.
The ZINC-12k dataset (Dwivedi et al., 2023), which contains 12 0001200012\,00012 000 molecules, is a subset of the larger ZINC dataset (Sterling & Irwin, 2015), which contains 250 000250000250\,000250 000 molecules. In ZINC-12k, each node has only one feature, which encodes the type of atom. The task is to regress constrained solubility, defined as the water-octanol partition coefficient ($logP$) minus the synthetic accessibility score ($SAS$) and the number of cycles with more than six atoms. Constrained solubility has implications in drug design, and thus has been used as the target for molecular graph generation models, as in (Jin et al., 2018) and (Kusner et al., 2017).
@@
We thus opt to use bootstrapping, sampling 10 0001000010\,00010 000 times with a confidence interval of $99\%$, yielding the results shown in Table 14. As Rings was developed from first principles, we apply the same classification schema to the relevance of information in each mode and the overall dataset evaluation.
@@
QM9 cf $<$ cg $<$ o $<$ eg $<$ rg $<$ rf uninformative unformative --- -
ZINC-12k o $<$ eg $<$ cg $<$ rg/cf $<$ rf informative informative +++++ +
@@
Interpretation of results. For QM9, performance-separability results suggest that neither the original mode nor its perturbations are particularly informative for the task. Notably, the original mode underperforms both $\operatorname{\operatorname{\varphi}_{cf}}$ and $\operatorname{\operatorname{\varphi}_{cg}}$, raising concerns about its suitability as a default configuration. Depending on the diversity of available modes, this may warrant realignment—or even deprecation—of QM9 in its current form. In contrast, ZINC-12k yields more encouraging results: The original mode consistently outperforms all perturbations, indicating that it captures task-relevant structure, and yielding a +++++ + overall classification. Interestingly, graph perturbations tend to outperform feature perturbations, suggesting that node features play a more critical role in this task than the underlying graph structure. For both datasets, a more exhaustive training routine—as employed in the main paper—is necessary to draw definitive conclusions regarding their placement within the regression taxonomy.
@@
As recommended by our reviewers, we further demonstrate that Rings also generalizes nicely to transformer architectures such as the GPS Transformer, introduced in (Rampásek et al., 2022). In contrast to GCN, GIN and GAT, GPS combines local message passing with global attention to overcome the expressivity and scalability limits of traditional GNNs. We compare the performance separability when considering 4 architectures across three datasets (NCI1, MUTAG, Proteins).
@@
MUTAG Present o/rg/cf/cg $>$ eg $>$ rf o/cf/rg/cg/eg/rf uninformative uninformative --- -
Absent o/rg/cf/cg $>$ eg $>$ rf o/cf/rg/cg/eg $>$ rf uninformative uninformative --- -
NCI1 Present o $>$ cg/cf $>$ rg $>$ eg/rf o $>$ cf/cg $>$ rg $>$ rf/eg informative informative +++++ +
Absent o $>$ cg/cf $>$ rg $>$ eg/rf o $>$ cf/cg $>$ rg $>$ rf $>$ eg informative informative +++++ +
Proteins Present cf $>$ cg $>$ o/rf/rg/eg eg/o/cg/rg $>$ rf/cf uninformative (un)informative --- -
Absent cf $>$ cg $>$ o/rf/rg/eg eg/o/cg/rg $>$ rf/cf uninformative (un)informative --- -
@@ D.3 Extended Mode Complementarity
Figure 9: Comparing levels of mode complementarity across different versions of the same dataset. We show the mean (dot) and $95$th percentile intervals (bars) of complementarity scores for the original version as well as $4$ deterministic perturbations (top) and $4$ randomized perturbations (bottom) of $7$ OGB datasets, computed with $t=1$ diffusion steps. Black horizontal lines indicate mean mode complementarities of the original dataset, and the silver horizontal line marks the $0.5$ threshold relevant for assessing mode diversity. Note that $\operatorname{\gamma}_{\text{eg}}=1-\operatorname{\gamma}_{\text{cg}}$ and $\operatorname{\gamma}_{\text{ef}}=1-\operatorname{\gamma}_{\text{cf}}$ by definition. Contrasting with the variation apparent in Figure 4, the OGB mol-$\ast$ datasets are remarkably similar in their mode-complementarity profiles.
@@ D.3.2 Mode Complementarity and Performance
Our intuition is that datasets with high complementarity (different and high information content across the two modes) should perform better. In line with this expectation, we observe a generally positive relationship between mode complementarity and performance when measured by both AUROC and accuracy, as visualized in Figure 5 and Figure 10, respectively. This correlation is further quantified in Table 18, where we symbolically categorize the observed correlations by creating five equal-width bins in the interval $[-1.0,1.0]$, i.e., $[-1.0,-0.6)$, $[-0.6,-0.2)$, $[-0.2,0.2)$, $[0.2,0.6)$, $[0.6,1.0]$. Again, this categorization only serves to communicate our main message (here: that the correlations are mostly positive and substantively robust to different ways of measurement), but the numerical values should be consulted for gaining detailed insights.
@@
DD -0.17 -0.31 -0.20 -0.27 -0.24 -0.16 $\circ$ - $\circ$ - - $\circ$ -
@@
Table 18: Mode complementarity and performance: Correlation statistics. We show the Pearson correlation between mode complementarity and test accuracy resp. AUROC for $13$ graph-learning datasets, taken over $5$ perturbations and $t\in[10]$ diffusion steps, based on our best-on-average models (as in Figure 3).

### v3

@@ D.2.4 Supplementary Extensibility Experiments
[…]
@@
Global (PyTorch) Seeds $\{$42, 7, 123, 56, 89}\}
Test/Train Split Seeds $\{$67, 23, 77, 88, 54}\}
Transform Seeds Random Modes: $\{$34, 12, 99, 45, 10}\}
Fixed Modes: $\{$34}\}
@@
Dataset descriptions. QM9 (Wu et al., 2018a) is a molecular dataset of 133 885133\,885 graphs. Its name originates from its focus on quantum mechanical properties and the makeup of the dataset: stable, organic molecules with nine or fewer heavy atoms. The PyG version has eleven node features, the first five being a one-hot encoding of the atom type, and the rest representing measures such as the hybridization state and the number of hydrogen atoms. Nineteen regression tasks seek to predict a variety molecular properties. We focus on task 0 (as defined by PyG), regressing the dipole moment, which describes the spatial distribution of electrons.
The ZINC-12k dataset (Dwivedi et al., 2023), which contains 12 00012\,000 molecules, is a subset of the larger ZINC dataset (Sterling & Irwin, 2015), which contains 250 000250\,000 molecules. In ZINC-12k, each node has only one feature, which encodes the type of atom. The task is to regress constrained solubility, defined as the water-octanol partition coefficient ($logP$) minus the synthetic accessibility score ($SAS$) and the number of cycles with more than six atoms. Constrained solubility has implications in drug design, and thus has been used as the target for molecular graph generation models, as in (Jin et al., 2018) and (Kusner et al., 2017).
@@
We thus opt to use bootstrapping, sampling 10 00010\,000 times with a confidence interval of $99\%$, yielding the results shown in Table 14. As Rings was developed from first principles, we apply the same classification schema to the relevance of information in each mode and the overall dataset evaluation.
@@
QM9 cf << cg << o << eg << rg << rf uninformative unformative --
ZINC-12k o << eg << cg << rg/cf << rf informative informative ++++
@@
Interpretation of results. For QM9, performance-separability results suggest that neither the original mode nor its perturbations are particularly informative for the task. Notably, the original mode underperforms both $\operatorname{\operatorname{\varphi}_{cf}}$ and $\operatorname{\operatorname{\varphi}_{cg}}$, raising concerns about its suitability as a default configuration. Depending on the diversity of available modes, this may warrant realignment—or even deprecation—of QM9 in its current form. In contrast, ZINC-12k yields more encouraging results: The original mode consistently outperforms all perturbations, indicating that it captures task-relevant structure, and yielding a ++++ overall classification. Interestingly, graph perturbations tend to outperform feature perturbations, suggesting that node features play a more critical role in this task than the underlying graph structure. For both datasets, a more exhaustive training routine—as employed in the main paper—is necessary to draw definitive conclusions regarding their placement within the regression taxonomy.
@@
As recommended by our reviewers, we further demonstrate that Rings also generalizes nicely to transformer architectures such as the GPS Transformer, introduced by Rampásek et al. (2022). In contrast to GCN, GIN and GAT, GPS combines local message passing with global attention to overcome the expressivity and scalability limits of traditional GNNs. We compare the performance separability when considering 4 architectures across three datasets (NCI1, MUTAG, Proteins). Note that other transformer architectures, such as the Graphformer architecture proposed by Yang et al. (2021) that we originally planned to include in our extended experiments, are tailored to scenarios with non-graph data, which lie beyond the scope of this work.
@@
MUTAG Present o/rg/cf/cg >> eg >> rf o/cf/rg/cg/eg/rf uninformative uninformative --
Absent o/rg/cf/cg >> eg >> rf o/cf/rg/cg/eg >> rf uninformative uninformative --
NCI1 Present o >> cg/cf >> rg >> eg/rf o >> cf/cg >> rg >> rf/eg informative informative ++++
Absent o >> cg/cf >> rg >> eg/rf o >> cf/cg >> rg >> rf >> eg informative informative ++++
Proteins Present cf >> cg >> o/rf/rg/eg eg/o/cg/rg >> rf/cf uninformative (un)informative --
Absent cf >> cg >> o/rf/rg/eg eg/o/cg/rg >> rf/cf uninformative (un)informative --
@@ D.3 Extended Mode Complementarity
Figure 9: Comparing levels of mode complementarity across different versions of the same dataset. We show the mean (dot) and $95$th percentile intervals (bars) of complementarity scores for the original version as well as 44 deterministic perturbations (top) and $4$ randomized perturbations (bottom) of 77 OGB datasets, computed with $t=1$ diffusion steps. Black horizontal lines indicate mean mode complementarities of the original dataset, and the silver horizontal line marks the $0.5$ threshold relevant for assessing mode diversity. Note that $\operatorname{\gamma}_{\text{eg}}=1-\operatorname{\gamma}_{\text{cg}}$ and $\operatorname{\gamma}_{\text{ef}}=1-\operatorname{\gamma}_{\text{cf}}$ by definition. Contrasting with the variation apparent in Figure 4, the OGB mol-$\ast$ datasets are remarkably similar in their mode-complementarity profiles.
@@ D.3.2 Mode Complementarity and Performance
Our intuition is that datasets with high complementarity (different and high information content across the two modes) should perform better. In line with this expectation, we observe a generally positive relationship between mode complementarity and performance when measured by both AUROC and accuracy, as visualized in Figure 5 and Figure 10, respectively. This correlation is further quantified in Table 18, where we symbolically categorize the observed correlations by creating five equal-width bins in the interval [1.0,1.0][-1.0,1.0], i.e., [1.0,0.6)[-1.0,-0.6), [0.6,0.2)[-0.6,-0.2), [0.2,0.2)[-0.2,0.2), $[0.2,0.6)$, $[0.6,1.0]$. Again, this categorization only serves to communicate our main message (here: that the correlations are mostly positive and substantively robust to different ways of measurement), but the numerical values should be consulted for gaining detailed insights.
@@
$D$ -0.17 -0.31 -0.20 -0.27 -0.24 -0.16 $\circ$ - $\circ$ - - $\circ$ -
@@
Table 18: Mode complementarity and performance: Correlation statistics. We show the Pearson correlation between mode complementarity and test accuracy resp. AUROC for $13$ graph-learning datasets, taken over 55 perturbations and $t\in[10]$ diffusion steps, based on our best-on-average models (as in Figure 3).